3x-5 = 7x+4
------ -------
6 4
Can you help me solve this equation step by step

Answer by stanbon(75887)

3x-5 = 7x+4
------ -------
6 4
========================
Cross-multiply:
4(3x-5) = 6(7x+4)
-----
12x - 20 = 42x + 24
-----
30x = -44
-----
x = -22/15
------------
Cheers,
Stan H.
=============